OpenStudy (brydenwright):

Eliminate the parameter. x = 4 cos t, y = 4 sin t

OpenStudy (solomonzelman):

\(\cos^2t+\sin^2t=1\) Right?

OpenStudy (brydenwright):

Yeah

OpenStudy (solomonzelman):

OK, so for example, if I had: x = r cos t, y = r sin t then, I would eliminate the parameters like this: \(x^2+y^2=(r \cos t)^2+(r\sin t)^2\) \(x^2+y^2=r^2\cos^2t+r^2\sin^2t\) \(x^2+y^2=r^2(\cos^2t+\sin^2t)\) \(x^2+y^2=r^2\)

OpenStudy (brydenwright):

Oh, wow. That makes sense. Thanks for the help!

OpenStudy (solomonzelman):

So, you know that \(y=r\sin t\) and \(x=r\cos t\) just give the equation of a circle with radius of \(r\), and centered at the origin.

OpenStudy (solomonzelman):

So, your answer would be (if you don't mind) ?

OpenStudy (sshayer):

\[\cos t=\frac{ x }{ 4 },\sin t=\frac{ y }{ 4 }\] \[\cos ^2t+\sin ^2t=1\]